# RMT-KD: Random Matrix Theoretic Causal Knowledge Distillation

### v1

@@ I Introduction and Prior Works
Recent advances in natural language processing (NLP) and computer vision rely on increasingly large deep learning models such as BERT [1] and convolutional networks such as ResNet [2]. While highly accurate, these models incur substantial inference latency, memory footprint, and carbon emissions [3]. Classical compression methods include knowledge distillation (KD), which transfers predictions from a frozen teacher to a smaller student [4], with refinements such as DistilBERT [5] and progressive shrinking [6]; pruning, which removes unimportant weights or channels [7]; and low-rank factorisation, which compresses convolutional filters [8] and adaptation matrices in large language models (LLMs) [9]. Yet these techniques often depend on heuristic thresholds, yield hardware-unfriendly sparsity, or lack a principled statistical rule.
In this paper, we ask: Is there a principled statistical basis for model reduction? We argue that Random Matrix Theory (RMT) provides such a foundation [10]. In high-dimensional settings, the eigenvalue spectrum of activation covariances typically separates into a bulk that reflects random noise and a small number of outliers that capture structured, causal features [11, 12]. This separation, formalized by the Marchenko–Pastur (MP) law [13], offers a rigorous criterion for identifying informative directions.
@@ II RMT-KD: Methodology
We propose an iterative self-distillation method that compresses neural networks by progressively reducing their width under RMT guidance. Unlike standard knowledge distillation, where a student learns from a fixed teacher, our approach uses a single model that distills itself at intermediate checkpoints. Each iteration treats the current model as the teacher and its reduced counterpart as the student, trained with a combined loss:

### v3

@@ I Introduction and Prior Works
Recent advances in natural language processing (NLP) and computer vision rely on increasingly large deep learning models such as BERT [1] and convolutional networks such as ResNet [2]. While highly accurate, these models incur substantial inference latency, memory footprint, and carbon emissions [3]. Classical compression methods include knowledge distillation (KD), which transfers predictions from a frozen teacher to a smaller student [4], with refinements such as DistilBERT [5] and progressive shrinking [6]; pruning, which removes unimportant weights or channels [7]; and low-rank factorisation, which compresses convolutional filters [8] and adaptation matrices in large language models [9]. Yet these methods depend on heuristic thresholds, yield hardware-unfriendly sparsity, or lack a principled statistical rule.
In this paper, we ask: Is there a principled statistical basis for model reduction? We argue that Random Matrix Theory (RMT) provides such a foundation [10]. In high-dimensional settings, the eigenvalue spectrum of activation covariances typically separates into a bulk that reflects random noise and a small number of outliers that capture structured, causal features [11, 12]. This separation, formalized by the Marchenko–Pastur (MP) law [13], offers a rigorous criterion for identifying informative directions, as shown in [14].
@@ II RMT-KD: Methodology
We propose an iterative self-distillation method that compresses neural networks by progressively reducing their width under RMT guidance. Unlike standard knowledge distillation, where a student learns from a fixed teacher, our approach uses a single model that distills itself at intermediate checkpoints. Each iteration treats the current model as teacher and its reduced counterpart as student, trained with a combined loss: